# Navigating Distribution Shifts in Medical Image Analysis: A Survey

### v1

@@ V Federated Learning
Similar to the Joint Training category, Federated Learning methods can also be classified based on the degree of data labeling. However, unlike Joint Training, which primarily focuses on the label availability of target data, Federated Learning treats both source and target data as clients that play similar roles. Each client trains a local model that contributes to the quality of the expected global model. Therefore, in this section, our primary concern regarding label availability extends to all clients. Based on this, Federated Learning for MedIA under distribution shifts can be divided into Supervised and Semi-supervised Federated Learning. Besides, it is worth noting that, given that medical settings often feature well-characterized source datasets, the need for Unsupervised Federated Learning approaches [121, 122] is generally minimal.
V-A Superivsed Federated Learning
@@
Early research in the medical field utilizing the FedAvg [25] algorithm targeted a wide range of medical tasks, from Brain Tumor Segmentation [123] to the detection of COVID-19 lung abnormalities [124, 125], MRI Reconstruction [126], Diabetic Retinopathy Classification of OCT Data [127], and Breast Cancer Histopathological Image Classification [128]. These initial applications laid the foundation for using Federated Learning to process sensitive medical data across distributed datasets while maintaining privacy. Further studies explored the influence of factors like the number of healthcare providers, dataset size, communication strategies, and architecture types on FL performance in medical contexts [129, 130]. As the field progressed, benchmarks were established to assess the effectiveness of various FL algorithms in managing data heterogeneity [131] across diverse medical datasets. Moving beyond the basic FedAvg paradigm, current advancements have focused on addressing issues like Data Heterogeneity and Client Drift [132, 133], which arise from non-IID data distributions among clients. These shifts can significantly affect model performance, prompting researchers to develop strategies for the accuracy of FL models in healthcare settings.
V-A1 Data Management
@@
Data Augmentation. In Federated Learning for MedIA under distribution shifts, cross-client data augmentation plays a crucial role in managing the inherent diversity and imbalance of data across different clients. This strategy is designed to enhance the uniformity of feature representations across participating clients, thus improving the overall robustness and accuracy of the federated model. Techniques such as Fourier transform-based methods [134] are particularly effective, as they allow for the sharing and interpolation of frequency domain information among clients, promoting a more consistent feature representation across varied datasets. Specifically, HarmoFL [135] leverages frequency information to unify amplitude components across clients, which aids in maintaining consistent low-level visual features. Other systematic augmentation techniques [136, 137, 138, 139] explore various augmentation strategies to combat data diversity and imbalance. These techniques vary in their approaches but collectively contribute to a more equitable and effective training process, enhancing the ability of federated models to generalize across diverse environments and data conditions.
V-A2 Model Design
@@
Novel Architecture Design. Some strategies specifically address distribution shifts by developing tailored model architectures. For instance, SU-Net [140] enhances standard U-Net with inception modules and dense blocks to manage multi-scale challenges effectively. Similarly, FedDAvT [141] leverages Transformer architecture to facilitate domain adaptation for Alzheimer’s disease diagnostics. Adversarial and generative networks are introduced to refine federated learning, focusing on aligning or adapting feature spaces across different clients [142, 143].
V-A3 Optimization Strategy
@@
Metric Learning. Several methods utilize metric learning to enhance consistency between different clients in federated settings. For instance, FedIIC [144] implements two-level contrastive learning to optimize both intra- and inter-client feature consistency, ensuring uniformity in the learned representations. FedCL [145] focuses on reducing the feature distance between successive local and global models, which helps stabilize the training process. Similarly, FedDP [146] improves model uniformity across clients by penalizing inconsistencies during the learning phase. Additionally, LC-Fed [147] employs contrastive site embedding and makes prediction-level adjustments to enhance personalization.
@@
Novel Training Strategies. Novel Training Strategies are being explored to enhance the efficacy and adaptability of models. For example, FedSM [148] optimizes model selection based on data distribution during inference, while FedCross [149] employs a unique approach that involves sequential training without the need for model aggregation. Additionally, strategies such as the Dropout, Mixture of Experts and Split Learning have been introduced to improve model effectiveness [150, 151, 152, 153]. These innovative methods collectively contribute to more secure and resilient model training and deployment.
@@
Aggregation Weight Calibration. In Federated Learning for MedIA, aggregation weight calibration is a sophisticated optimization strategy that refines how global model updates are weighted, taking into account more than data volume. This method involves adjusting the influence of each client’s local update on the global model by considering factors such as the stage of training, client performance, and similarity between client models and the global model. For example, [154, 155] highlight strategies where weights are calibrated based on the training progress and the performance metrics of clients. Additionally, the similarity-based approach [156, 157, 158, 159] assesses how closely aligned each client’s data distribution or model parameters are with the global model. This alignment influences their weights during aggregation, promoting updates that are more representative of the overall data characteristics. Moreover, FedAWA [160] introduces an innovative twist by employing reinforcement learning to dynamically adjust client weights. This system continually learns and updates based on data distribution and feedback from client performance, optimizing the aggregation process to ensure the global model remains robust and accurate across varying conditions.
@@
Parameter Calibration. Parameter calibration also plays a crucial role, specifically for addressing the conflict between the local and global models. It involves strategically adjusting model parameters to ensure that the collective learning benefits all participating clients. Efforts include rescaling local parameters [161] and mixing local and global gradients [162] to enhance model convergence and stability. [163] proposes a Deputy-Enhanced Transfer strategy at the client site. It firstly leverages a deputy model to receive aggregated parameters from the server, and then smoothly transfers the global knowledge to the personalized local model. Some other strategies emphasize fairness, such as those aiming to equalize training loss by adjusting the model parameters such that all hospitals have a similar training loss [164]. This approach ensures that no single client’s data disproportionately influences the model, thus maintaining uniformity in model performance regardless of the data source.
V-B Semi-supervised Federated Learning
@@
V-B1 Data Management
@@
Pseudo-labeling. Several innovative approaches have been developed to enhance the utility of pseudo-labeling in federated settings. [165] introduces a novel method that integrates prototype-based pseudo-labeling with contrastive learning, a technique also employed by [166]. Additionally, [167] enhances pseudo-label generation by incorporating a self-supervised rotation loss, which provides consistent regularization across unlabeled datasets. Further, [168] improves the connection between labeled and unlabeled data by aligning disease relationships across clients, effectively compensating for the lack of task-specific knowledge in unlabeled clients and enhancing the extraction of discriminative information from unlabeled samples.
V-B2 Model Design
@@
Transformer-based Architecture. Transformer offers a robust framework for leveraging both labeled and unlabeled data within a single client. For example, [169] exemplifies a specialized approach where a self-supervised learning framework is implemented using Transformer architectures. This method starts with masked image modeling, a self-supervised task that trains the model to predict the portions of images that are intentionally obscured. This phase harnesses the abundant unlabeled data, allowing the model to learn rich, generalized features without requiring too many explicit labels.
V-B3 Optimization Strategy
@@
Advanced Optimization Strategies in SSFL also address the dual challenges of data scarcity and distribution heterogeneity. One innovative approach is the Federated Drift Mitigation (FedDM) framework [170], which achieves robust gradient aggregation by resolving conflicts between gradients at different network layers, as guided by the historical gradients of the global model. Another strategic implementation is FedCy [171], designed for surgical phase recognition. This method integrates dual training objectives: it applies consistency learning to exploit the temporal and spatial consistencies in the unlabeled data, alongside contrastive learning techniques to enrich the learning from sparsely labeled data.
VI Fine-tuning
@@
Fine-tuning plays a vital role in enhancing the adaptability and performance of pre-trained models across a wide range of applications. This process involves adjusting a model that has been pre-trained on a large, generic (source) dataset to perform effectively on a different, often smaller and more specialized (target) dataset. In medical scenarios, Fine-tuning proves particularly effective when privacy concerns preclude open data sharing, and synchronous collaborations among different healthcare institutions are impractical or excessively costly. This strategy enables medical institutions to leverage pre-existing models and adapt them with minimal data exchange, effectively addressing privacy and collaboration constraints in MedIA. Based on the availability of labeled data on the target domain, fine-tuning methods are classified into supervised and unsupervised approaches. As we move from supervised to unsupervised settings, the complexity increases but so does the significance of the application, offering broader adaptability to real-world challenges where labeled data are limited.
VI-A Supervised Fine-tuning
@@
Supervised Fine-tuning stands out as a potent method for enhancing diagnostic accuracy in MedIA. This technique primarily involves applying specific pre-trained networks, such as VGG [172] and AlexNet [173], initially trained on general images like ImageNet [174], to more specialized medical imaging tasks. Research exemplified by studies [175, 176, 177, 178] demonstrates how these models transition to applications in medical imaging, including tumor classification and chest X-ray analysis, leveraging their capability to generalize features across diverse visual domains for precise medical diagnostics. Fine-tuning these networks often requires minimal adaptation design, making it a straightforward approach to boost performance in medical tasks. Notable successes also include adapting networks for Alzheimer’s diagnosis [26] and employing the Med3D network for detailed lung segmentation and nodule classification [179].
VI-A1 Model Design
@@
Novel Strategies and Structures. Beyond simply evaluating on different pre-trained network architectures, some research have focused on novel strategies and structures for rapid and accurate domain adaptation, while preserving existing knowledge. [180] introduces ContextNets, a memory-augmented network for seamless domain adaptation in semantic segmentation without the need for extensive retraining. In contrast, [181] employs Elastic Weight Consolidation to maintain performance by encoding information from previous tasks, without extra data storage. Furthermore, [182] optimizes batch normalization to swiftly adjust to new domains while maintaining shared convolutional layers across all domains.
VI-B Unsupervised Fine-tuning
@@
Unsupervised Fine-tuning in MedIA is an innovative response to the constraints of traditional supervised fine-tuning that rely heavily on labeled target datasets which are often unavailable in healthcare scenarios. This approach, crucial in healthcare where rapid adaptation is required to varying patient data, is characterized by two primary branches: Source-Free Domain Adaptation (SFDA) [183] and Test-time Adaptation (TTA) [184]. Both are designed to adapt models dynamically to new and changing conditions without the need for source data at the time of inference, thus directly addressing the challenges of data privacy. SFDA achieves this by transferring knowledge learned during training and applying it to new test samples through adaptive modules or auxiliary self-supervised tasks, such as rotation prediction. This allows the model to train on the target distribution for multiple epochs before making predictions, providing a proactive adaptation approach. On the other hand, TTA takes on a more challenging task by requiring the model to adapt in real-time to a continuous stream of test data, making no prior adjustments during the training phase. This method is model-agnostic and focuses on immediate, on-the-fly adjustments to effectively process and respond to incoming data. Both strategies share the common goal of enabling efficient model adaptation in unsupervised settings, ensuring that medical diagnostics remain robust and accurate even when faced with data that significantly deviates from previously seen examples.
VI-B1 Data Management
@@
Pseudo-labeling. An intuitive solution for SFDA/TTA is to use the source model to generate pseudo labels [185, 186, 187, 188] for the target domain data and thus convert the problem into a supervised one on the target domain. However, these pseudo labels often contain noise due to domain discrepancies, making it essential to refine them for accuracy. Techniques include adaptive pseudo-labeling which uses dual-classifiers to enhance label confidence in [189], denoised pseudo-labeling with uncertainty and prototype distance estimation for precise segmentation [190], and employing shape compactness metrics for label reweighting [191]. Additionally, in [192], a system integrates an image quality assessor and an irregular structure detector is developed to select optimal pseudo-labels for training. [193] uses the greatest union mask of multiple predictions to generate proxy labels for model fine-tuning, while [188] selects low-entropy pixels as reliable labels and applies contrastive learning to tighten the target feature distribution.
@@
Image Generation. Image generation techniques facilitate the adaptation of models to new domains by enriching the dataset with varied and representative examples. For example, [194] utilizes basic image augmentation combined with causal interventions to generate diverse datasets that ensure consistent predictions and the elimination of confounding factors. Similarly, [195] employs patch-wise processing augmented with a Transformer structure to enhance data variability effectively, while [196] proposes the first learnable test-time augmentation policy that dynamically selects most effective augmentation techniques. This adaptability allows for optimal model performance even under varying operational conditions. Moreover, some strategies focus on transforming the style of data between the source and target domains to better align the characteristics of the target data with the learned source domain model. For instance, [197] applies autoencoders to adjust test images to resemble source-domain images more closely, enhancing the model’s applicability to new data. Additionally, [198] and [191] explore generative techniques. The former uses a class-conditional generative adversarial network to create target-style data from random noise, while the latter leverages Fourier transformation to generate source-like images through a style-mining generator. [199] further innovates by learning a domain-aware prompt that modifies target inputs to better match the source domain style, facilitating smoother domain adaptation.
VI-B2 Model Design

### v2

@@ 6. Federated Learning
Similar to the Joint Training category, Federated Learning methods can also be classified based on the degree of data labeling. However, unlike Joint Training, which primarily focuses on the label availability of target data, Federated Learning treats both source and target data as clients that play similar roles. Each client trains a local model that contributes to the quality of the expected global model. Therefore, in this section, our primary concern regarding label availability extends to all clients. Based on this, Federated Learning for MedIA under distribution shifts can be divided into Supervised and Semi-supervised Federated Learning. Besides, it is worth noting that, given that medical settings often feature well-characterized source datasets, the need for Unsupervised Federated Learning approaches (Lubana et al., 2022; Kim et al., 2023) is generally minimal.
6.1. Superivsed Federated Learning
@@
Early research in the medical field utilizing the FedAvg (McMahan et al., 2017) algorithm targeted a wide range of medical tasks, from Brain Tumor Segmentation (Sheller et al., 2019) to the detection of COVID-19 lung abnormalities (Dou et al., 2021; Feki et al., 2021), MRI Reconstruction (Elmas et al., 2022), Diabetic Retinopathy Classification of OCT Data (Lo et al., 2021), and Breast Cancer Histopathological Image Classification (Li et al., 2022d). These initial applications laid the foundation for using Federated Learning to process sensitive medical data across distributed datasets while maintaining privacy. Further studies explored the influence of factors like the number of healthcare providers, dataset size, communication strategies, and architecture types on FL performance in medical contexts (Adnan et al., 2022; Tedeschini et al., 2022). As the field progressed, benchmarks were established to assess the effectiveness of various FL algorithms in managing data heterogeneity (Yang et al., 2022c) across diverse medical datasets. Moving beyond the basic FedAvg paradigm, current advancements have focused on addressing issues like Data Heterogeneity and Client Drift (Li et al., 2020b; Collins et al., 2021), which arise from non-IID data distributions among clients. These shifts can significantly affect model performance, prompting researchers to develop strategies for the accuracy of FL models in healthcare settings.
6.1.1. Data Management
6.1.1.1 Data Augmentation.
@@ 6.1.1. Data Management
In Federated Learning for MedIA under distribution shifts, cross-client data augmentation plays a crucial role in managing the inherent diversity and imbalance of data across different clients. This strategy is designed to enhance the uniformity of feature representations across participating clients, thus improving the overall robustness and accuracy of the federated model. Techniques such as Fourier transform-based methods (Liao et al., 2022) are particularly effective, as they allow for the sharing and interpolation of frequency domain information among clients, promoting a more consistent feature representation across varied datasets. Specifically, HarmoFL (Jiang et al., 2022b) leverages frequency information to unify amplitude components across clients, which aids in maintaining consistent low-level visual features. Other systematic augmentation techniques (Linardos et al., 2022; Cetinkaya et al., 2021; Yan et al., 2020; Manthe et al., 2024) explore various augmentation strategies to combat data diversity and imbalance. These techniques vary in their approaches but collectively contribute to a more equitable and effective training process, enhancing the ability of federated models to generalize across diverse environments and data conditions.
6.1.2. Model Design
6.1.2.1 Novel Architecture Design.
@@ 6.1.2. Model Design
Some strategies specifically address distribution shifts by developing tailored model architectures. For instance, SU-Net (Yi et al., 2020) enhances standard U-Net with inception modules and dense blocks to manage multi-scale challenges effectively. Similarly, FedDAvT (Lei et al., 2023) leverages Transformer architecture to facilitate domain adaptation for Alzheimer’s disease diagnostics. Adversarial and generative networks are introduced to refine federated learning, focusing on aligning or adapting feature spaces across different clients (Guo et al., 2021; Dalmaz et al., 2022).
6.1.3. Optimization Strategy
6.1.3.1 Metric Learning.
@@ 6.1.3. Optimization Strategy
Several methods utilize metric learning to enhance consistency between different clients in federated settings. For instance, FedIIC (Wu et al., 2023c) implements two-level contrastive learning to optimize both intra- and inter-client feature consistency, ensuring uniformity in the learned representations. FedCL (Liu et al., 2023b) focuses on reducing the feature distance between successive local and global models, which helps stabilize the training process. Similarly, FedDP (Wang et al., 2023) improves model uniformity across clients by penalizing inconsistencies during the learning phase. Additionally, LC-Fed (Wang et al., 2022a) employs contrastive site embedding and makes prediction-level adjustments to enhance personalization.
6.1.3.2 Novel Training Strategies.
@@
Novel Training Strategies are being explored to enhance the efficacy and adaptability of models. For example, FedSM (Xu et al., 2022a) optimizes model selection based on data distribution during inference, while FedCross (Xu et al., 2024) employs a unique approach that involves sequential training without the need for model aggregation. Additionally, strategies such as the Dropout, Mixture of Experts and Split Learning have been introduced to improve model effectiveness (Gunesli et al., 2023; Ding et al., 2023; Li et al., 2020a; Zhang et al., 2022b). These innovative methods collectively contribute to more secure and resilient model training and deployment.
6.1.3.3 Aggregation Weight Calibration.
@@
In Federated Learning for MedIA, aggregation weight calibration is a sophisticated optimization strategy that refines how global model updates are weighted, taking into account more than data volume. This method involves adjusting the influence of each client’s local update on the global model by considering factors such as the stage of training, client performance, and similarity between client models and the global model. For example, (Otálora et al., 2022; Shen et al., 2021) highlight strategies where weights are calibrated based on the training progress and the performance metrics of clients. Additionally, the similarity-based approach (Deng et al., 2023; Khan et al., 2022; Yeganeh et al., 2020; Zhang et al., 2023a) assesses how closely aligned each client’s data distribution or model parameters are with the global model. This alignment influences their weights during aggregation, promoting updates that are more representative of the overall data characteristics. Moreover, FedAWA (Yue et al., 2023) introduces an innovative twist by employing reinforcement learning to dynamically adjust client weights. This system continually learns and updates based on data distribution and feedback from client performance, optimizing the aggregation process to ensure the global model remains robust and accurate across varying conditions.
6.1.3.4 Parameter Calibration.
@@
Parameter calibration also plays a crucial role, specifically for addressing the conflict between the local and global models. It involves strategically adjusting model parameters to ensure that the collective learning benefits all participating clients. Efforts include rescaling local parameters (Yin et al., 2021) and mixing local and global gradients (Jiang et al., 2023) to enhance model convergence and stability. (Chen et al., 2021d) proposes a Deputy-Enhanced Transfer strategy at the client site. It firstly leverages a deputy model to receive aggregated parameters from the server, and then smoothly transfers the global knowledge to the personalized local model. Some other strategies emphasize fairness, such as those aiming to equalize training loss by adjusting the model parameters such that all hospitals have a similar training loss (Hosseini et al., 2023). This approach ensures that no single client’s data disproportionately influences the model, thus maintaining uniformity in model performance regardless of the data source.
6.2. Semi-supervised Federated Learning
@@
6.2.1. Data Management
6.2.1.1 Pseudo-labeling.
@@ 6.2.1. Data Management
Several innovative approaches have been developed to enhance the utility of pseudo-labeling in federated settings. (Wu et al., 2023d) introduces a novel method that integrates prototype-based pseudo-labeling with contrastive learning, a technique also employed by (Qiu et al., 2023). Additionally, (Liu et al., 2023a) enhances pseudo-label generation by incorporating a self-supervised rotation loss, which provides consistent regularization across unlabeled datasets. Further, (Liu et al., 2021h) improves the connection between labeled and unlabeled data by aligning disease relationships across clients, effectively compensating for the lack of task-specific knowledge in unlabeled clients and enhancing the extraction of discriminative information from unlabeled samples.
6.2.2. Model Design
6.2.2.1 Transformer-based Architecture.
@@ 6.2.2. Model Design
Transformer offers a robust framework for leveraging both labeled and unlabeled data within a single client. For example, (Yan et al., 2023) exemplifies a specialized approach where a self-supervised learning framework is implemented using Transformer architectures. This method starts with masked image modeling, a self-supervised task that trains the model to predict the portions of images that are intentionally obscured. This phase harnesses the abundant unlabeled data, allowing the model to learn rich, generalized features without requiring too many explicit labels.
6.2.3. Optimization Strategy
6.2.3.1 Advanced Optimization Strategies
@@
in SSFL also address the dual challenges of data scarcity and distribution heterogeneity. One innovative approach is the Federated Drift Mitigation (FedDM) framework (Zhu et al., 2023b), which achieves robust gradient aggregation by resolving conflicts between gradients at different network layers, as guided by the historical gradients of the global model. Another strategic implementation is FedCy (Kassem et al., 2022), designed for surgical phase recognition. This method integrates dual training objectives: it applies consistency learning to exploit the temporal and spatial consistencies in the unlabeled data, alongside contrastive learning techniques to enrich the learning from sparsely labeled data.
Figure 5. Fine-tuning in MedIA: It adapts pre-trained models to specialized datasets, particularly when privacy concerns limit data sharing and direct collaborations are impractical. This approach enables healthcare institutions to leverage the valuable prior knowledge with minimal data exchange.
7. Fine-tuning
@@
Fine-tuning plays a vital role in enhancing the adaptability and performance of pre-trained models across a wide range of applications. This process involves adjusting a model that has been pre-trained on a large, generic (source) dataset to perform effectively on a different, often smaller and more specialized (target) dataset. In medical scenarios, Fine-tuning proves particularly effective when privacy concerns preclude open data sharing, and synchronous collaborations among different healthcare institutions are impractical or excessively costly. This strategy enables medical institutions to leverage pre-existing models and adapt them with minimal data exchange, effectively addressing privacy and collaboration constraints in MedIA, as illustrated in Fig. 5. Based on the availability of labeled data on the target domain, fine-tuning methods are classified into supervised and unsupervised approaches. As we move from supervised to unsupervised settings, the complexity increases but so does the significance of the application, offering broader adaptability to real-world challenges where labeled data are limited.
7.1. Supervised Fine-tuning
@@
Supervised Fine-tuning stands out as a potent method for enhancing diagnostic accuracy in MedIA. This technique primarily involves applying specific pre-trained networks, such as VGG (Simonyan and Zisserman, 2014) and AlexNet (Krizhevsky et al., 2012), initially trained on general images like ImageNet (Deng et al., 2009), to more specialized medical imaging tasks. Research exemplified by studies (Khan et al., 2019; Chouhan et al., 2020; Abbas et al., 2020; Swati et al., 2019) demonstrates how these models transition to applications in medical imaging, including tumor classification and chest X-ray analysis, leveraging their capability to generalize features across diverse visual domains for precise medical diagnostics. Fine-tuning these networks often requires minimal adaptation design, making it a straightforward approach to boost performance in medical tasks. Notable successes also include adapting networks for Alzheimer’s diagnosis (Hosseini-Asl et al., 2016) and employing the Med3D network for detailed lung segmentation and nodule classification (Chen et al., 2019b).
7.1.1. Model Design
7.1.1.1 Novel Strategies and Structures.
@@ 7.1.1. Model Design
Beyond simply evaluating on different pre-trained network architectures, some research have focused on novel strategies and structures for rapid and accurate domain adaptation, while preserving existing knowledge. (Venkataramani et al., 2019) introduces ContextNets, a memory-augmented network for seamless domain adaptation in semantic segmentation without the need for extensive retraining. In contrast, (Lenga et al., 2020b) employs Elastic Weight Consolidation to maintain performance by encoding information from previous tasks, without extra data storage. Furthermore, (Karani et al., 2018) optimizes batch normalization to swiftly adjust to new domains while maintaining shared convolutional layers across all domains.
7.2. Unsupervised Fine-tuning
@@
Unsupervised Fine-tuning in MedIA is an innovative response to the constraints of traditional supervised fine-tuning that rely heavily on labeled target datasets which are often unavailable in healthcare scenarios. This approach, crucial in healthcare where rapid adaptation is required to varying patient data, is characterized by two primary branches: Source-Free Domain Adaptation (SFDA) (Chidlovskii et al., 2016) and Test-time Adaptation (TTA) (Wang et al., 2021). Both are designed to adapt models dynamically to new and changing conditions without the need for source data at the time of inference, thus directly addressing the challenges of data privacy. SFDA achieves this by transferring knowledge learned during training and applying it to new test samples through adaptive modules or auxiliary self-supervised tasks, such as rotation prediction. This allows the model to train on the target distribution for multiple epochs before making predictions, providing a proactive adaptation approach. On the other hand, TTA takes on a more challenging task by requiring the model to adapt in real-time to a continuous stream of test data, making no prior adjustments during the training phase. This method is model-agnostic and focuses on immediate, on-the-fly adjustments to effectively process and respond to incoming data. Both strategies share the common goal of enabling efficient model adaptation in unsupervised settings, ensuring that medical diagnostics remain robust and accurate even when faced with data that significantly deviates from previously seen examples.
7.2.1. Data Management
7.2.1.1 Pseudo-labeling.
@@ 7.2.1. Data Management
An intuitive solution for SFDA/TTA is to use the source model to generate pseudo labels (Wu et al., 2023b; Cheng et al., 2023; Zhou et al., 2023; Yu et al., 2023) for the target domain data and thus convert the problem into a supervised one on the target domain. However, these pseudo labels often contain noise due to domain discrepancies, making it essential to refine them for accuracy. Techniques include adaptive pseudo-labeling which uses dual-classifiers to enhance label confidence in (Li et al., 2022b), denoised pseudo-labeling with uncertainty and prototype distance estimation for precise segmentation (Chen et al., 2021c), and employing shape compactness metrics for label reweighting (Yang et al., 2022b). Additionally, in (Li et al., 2023b), a system integrates an image quality assessor and an irregular structure detector is developed to select optimal pseudo-labels for training. (Huang et al., 2023) uses the greatest union mask of multiple predictions to generate proxy labels for model fine-tuning, while (Yu et al., 2023) selects low-entropy pixels as reliable labels and applies contrastive learning to tighten the target feature distribution.
7.2.1.2 Image Generation.
@@
Image generation techniques facilitate the adaptation of models to new domains by enriching the dataset with varied and representative examples. For example, (Qiu, 2023) utilizes basic image augmentation combined with causal interventions to generate diverse datasets that ensure consistent predictions and the elimination of confounding factors. Similarly, (Ye et al., 2022) employs patch-wise processing augmented with a Transformer structure to enhance data variability effectively, while (Tomar et al., 2022b) proposes the first learnable test-time augmentation policy that dynamically selects most effective augmentation techniques. This adaptability allows for optimal model performance even under varying operational conditions. Moreover, some strategies focus on transforming the style of data between the source and target domains to better align the characteristics of the target data with the learned source domain model. For instance, (He et al., 2021) applies autoencoders to adjust test images to resemble source images more closely, enhancing the model’s applicability to new data. Additionally, (Zhou et al., 2022b) and (Yang et al., 2022b) explore generative techniques. The former uses a class-conditional generative adversarial network to create target-style data from random noise, while the latter leverages Fourier transformation to generate source-like images through a style-mining generator. (Hu et al., 2022c) further innovates by learning a domain-aware prompt that modifies target inputs to better match the source domain style, facilitating smoother domain adaptation.
7.2.2. Model Design
7.2.2.1 Batch Normalization.